Eve shift 29-Jan-00


Subject: Eve shift 29-Jan-00
From: Vivian O'Dell (odell@fnal.gov)
Date: Fri Feb 04 2000 - 15:44:37 CST


                Evening SHIFT SUMMARY

                  29-JAN-2000

 SG on shift.

 Water in KTeV washrooms is fixed. Do not plan to use
 washrooms elsewhere.

 We got searched and secured at about 18.00. Got new
 key which opens key tree. It is stored in key box in the
 control room. Before using it we still MUST call MCR.

 At 20.15 we took pedestal run. We observed that although the constants
 were loaded onto database, fpedcal did not see them. Paging Ed and looking
 into database more carefuly we found that recently offline version of
 ADGN bank was installed with unlimited run range. Since fpedcal is "offline"
 job it picked up these constants ignoring all others. Fixed by changing
 validity range of the offline bank.

 At ~21.00 while testing various KAMI configurations RK observed some neutrons
 in UCAL and 3pi0 events with far downstream vertex distribution.
 We learned that beam stop was not closed and thus these events
 is the first real KAMI data.

 Beam stop was closed and we continued taking muon data at very low intensity.
 Run 15638 is a first mu-run written to tapes.

 23:35 -- Tev quench at C4. 45min estimate for beam.

 



This archive was generated by hypermail 2b27 : Fri Feb 04 2000 - 15:44:40 CST